OpenStudy (anonymous):

The sales totals at Linda's food store have increased exponentially over the months. Which of these best shows the sales in the first three months? -$1200 in the first month,$1260 in the second month, and $1323 in the third month -$1200 in the first month, $1250 in the second month, $1300 in the third month -$1200 in the first month, $1272 in the second month, $1344 in the third month -$1200 in the first month, $1285 in the second month, $1370 in the third month

OpenStudy (anonymous):

@nincompoop can you help me please?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

@TypnoEchosis

OpenStudy (anonymous):

@KoalaGurl1

OpenStudy (threshsupport):

i will

OpenStudy (threshsupport):

do you know what exponential growth is?

OpenStudy (anonymous):

the pattern of the numbers?...not exactly sure

OpenStudy (threshsupport):

growth whose rate becomes ever more rapid in proportion to the growing total number or size

OpenStudy (threshsupport):

thats the exact definition

OpenStudy (threshsupport):

well you are looking for the third month to be a bigger growth that the second one

OpenStudy (anonymous):

which all 4 are

OpenStudy (threshsupport):

no

OpenStudy (anonymous):

i got c wrong the first time

OpenStudy (threshsupport):

a is the only one that works

OpenStudy (threshsupport):

i think thats it

OpenStudy (threshsupport):

a goes up by 3

OpenStudy (threshsupport):

b had no growth

OpenStudy (threshsupport):

c had no growth

OpenStudy (threshsupport):

and d had no growth

OpenStudy (threshsupport):

i would go with A :D

OpenStudy (anonymous):

ok ty can you help me with some more?

OpenStudy (threshsupport):

sure

OpenStudy (anonymous):

The same amount of trash is dumped into a landfill every day. The function below shows the total number of tons of trash n in the landfill after x days: n = 2000x + 1000 What does the number 2000 represent? Amount of trash originally in the landfill Amount of trash added to the landfill every day Rate at which the total trash increases every day Rate at which the total trash is recycled every day

OpenStudy (anonymous):

Midge added some chlorine to the water in a pool. The chlorine evaporated at a fixed rate every week. The table below shows the amount of chlorine f(n), in ounces, that was left in the pool after n weeks: n f(n) 1 16 2 8 3 4 4 2 Which function best shows the relationship between n and f(n)? f(n) = 16(0.5)n−1 f(n) = 32(0.5)n−1 f(n) = 16(0.5)n f(n) = 32(0.5)n+1

OpenStudy (threshsupport):

Rate at which the total trash increases every day

OpenStudy (threshsupport):

f(n) = 16(0.5)n

OpenStudy (threshsupport):

what grade ru in

OpenStudy (anonymous):

how i thought x was the number of tons of trash...what does the 1000 stand for then?

OpenStudy (threshsupport):

base amount of trash

OpenStudy (threshsupport):

the y intercept

OpenStudy (threshsupport):

so Y=5x+7<--- this one

OpenStudy (threshsupport):

if you were graphing

OpenStudy (anonymous):

ty.....

OpenStudy (threshsupport):

any more?

OpenStudy (anonymous):

A ball is dropped from a certain height. The function below represents the height f(n), in feet, to which the ball bounces at the nth bounce: f(n) = 9(0.7)n What does the number 9 in the function represent?

OpenStudy (threshsupport):

the initial height

OpenStudy (threshsupport):

from where it was first dropped

OpenStudy (anonymous):

Number of times the ball was dropped. Height from which the ball was dropped. Height to which the ball will bounce at the nth bounce. Number of times the ball bounces at the nth bounce.

OpenStudy (threshsupport):

B

OpenStudy (anonymous):

sooo b?

OpenStudy (threshsupport):

yes

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

he table below shows the functions to represent the sale f(t), in thousands, of four companies in t years: Company A B C D Sale f(t) = 0.87t f(t) = 0.756t f(t) = 1.26t f(t) = 1.09t Which of the following companies showed an exponential growth in sale? Company A and Company B Company B and Company C Company A and Company D Company C and Company D

OpenStudy (anonymous):

i didnt know exactly what to do

OpenStudy (threshsupport):

D

OpenStudy (threshsupport):

you are looking for a number greater than one

OpenStudy (anonymous):

how? sorry just need to know so when i speak with my teacher

OpenStudy (threshsupport):

if it was exponential decay it would be the opposite

OpenStudy (anonymous):

soo it has to be above 1?

OpenStudy (threshsupport):

I'm writing what to tell your teacher for you

OpenStudy (anonymous):

An expression is shown below: square root of 18 plus square root of 2 Which statement is true about the expression? It is rational and equal to 3. It is rational and equal to 4. It is irrational and equal to 3 multiplied by the square root of 2. It is irrational and equal to 4 multiplied by the square root of 2.

OpenStudy (anonymous):

ok

OpenStudy (threshsupport):

companies C and D are growing at a rate of 126% every year and 109% each year. this means they are growing in sales

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

i have some more

OpenStudy (threshsupport):

not sure on that one

OpenStudy (threshsupport):

the wording doesn't make sense

OpenStudy (threshsupport):

its not a or b

OpenStudy (anonymous):

Henry is taking a medicine for a common cold. The table below shows the amount of medicine f(t), in mg, that was present in Henry's body after time t: t (hours) 1 2 3 4 5 f(t) (mg) 282 265.08 249.18 234.22 220.17 Greg was administered 200 mg of the same medicine. The amount of medicine in his body f(t) after time t is shown by the equation below: f(t) = 200(0.88)t Which statement best describes the rate at which Henry's and Greg's bodies eliminated the medicine? Henry's body eliminated the antibiotic faster than Greg's body. Henry's body eliminated the antibiotic at the same rate as Greg's body. Henry's body eliminated the antibiotic at half of the rate at which Greg's body eliminated the antibiotic. Henry's body eliminated the antibiotic at one-fourth of the rate at which Greg's body eliminated the antibiotic.

OpenStudy (threshsupport):

I'm trying to decide if its c or d

OpenStudy (threshsupport):

C on the medicine one

OpenStudy (anonymous):

ok how do i know that these answers are all correct?

OpenStudy (anonymous):

sorry dont wanna get them wrong

OpenStudy (threshsupport):

i learned this last year

OpenStudy (threshsupport):

im a sophmore

OpenStudy (anonymous):

oh ok

OpenStudy (threshsupport):

im 90% sure the ones i told you are right

OpenStudy (anonymous):

im 8th grade

OpenStudy (threshsupport):

nice.... high school sucks

OpenStudy (threshsupport):

do you do online school?

OpenStudy (threshsupport):

i switched after freshman year

OpenStudy (threshsupport):

i gtg tag me if you have any more questions..

OpenStudy (threshsupport):

ill be on tomorrow

OpenStudy (threshsupport):

bye

OpenStudy (anonymous):

ok ty